The Everton College store paid $1906 for an order of 48 calculators. The store paid $10 for each scientific calculator. The others, all graphic calculators, cost the store $56 each. How many of each type of calculator was ordered? The store ordered _______ scientific calculators and ________ graphing calculators.
Answer by richwmiller(17219)

s+g=48
10s+56g=1906
g=31 s=17
check
10*17+31*56=1906
170+1736=1906
ok